OpenStudy (anonymous):

Verify the identity. sin 4u = 2 sin 2u cos 2u and cos ( x + pi/2) = -sin x

OpenStudy (anonymous):

second one you can use the "addition angle" formula \[\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\]

OpenStudy (anonymous):

you get \[\cos(x+\frac{\pi}{2})=\cos(x)\cos(\frac{\pi}{2})-\sin(x)\sin(\frac{\pi}{2})\] \[=\cos(x)\times 0-\sin(x)\times 1=-\sin(x)\]

OpenStudy (anonymous):

would cos ( x + pi/2) = -sinx = -sin x=sin(x)

OpenStudy (anonymous):

so what is the first one

OpenStudy (anonymous):

and is that your final answer for the second one

OpenStudy (anonymous):

\[\sin(2x)=2\cos(x)\sin(x)\] replace \(x\) by \(2u\)

OpenStudy (anonymous):

and this is for the first one or the second one